find the exact value of the expression sin(pie/6)cos(pie/3)+sin(pie/3)cos(pie/6)
Found 2 solutions by checkley77, jim_thompson5910:
Answer by checkley77(12569)

sin(pie/6)cos(pie/3)+sin(pie/3)cos(pie/6)
.00913+.01827=.0274 ans.

Answer by jim_thompson5910(28476)



sin(pi/6)cos(pi/3)+sin(pi/3)cos(pi/6) = sin(pi/6)cos(pi/3)+cos(pi/6)sin(pi/3)
                                      = sin(pi/6+pi/3)
                                      = sin(pi/6+2pi/6)
                                      = sin((3pi)/6)
                                      = sin(pi/2)
                                      = 1



So
sin(pi/6)cos(pi/3)+sin(pi/3)cos(pi/6) = 1